The SI and CGS unit of energy are joule and erg respectively.How many ergs are equal to one joule?

Grade:12

3 Answers

faisal
6 years ago
The S.I Unit of energy are joules and E.R.G(smallest unit of energy) respective .how E.R.G are to 1 joule.
Mohd Mujtaba
6 years ago
The answer is-((1joule=10^7erg.))Or you can say ((1erg=10^-7joule.))Hope it will help you Thank and buy.
Rohit
6 years ago
Joule and erg are the units which can be also used to express energy. We can find the conversion factor by using dimensional analysis or by units. For direct calculation 1 Joule = 10^7 ergs